OpenStudy (anonymous):

What is the slope of the line that passes through the points (1, 7) and (–4, –8)? A. –3 B. -1/3 C. 1/3 D. 3

OpenStudy (bmk614):

\[\frac{ y1-y2 }{ x1-x2 } = \frac{ 7-(-8) }{ 1-(-4) } = \frac{ 15 }{ 5 } = 3\] So the answer is D.

OpenStudy (anonymous):

thanks

OpenStudy (bmk614):

\[\frac{ y2-y1 }{ x2-x1 }=\frac{ -8-7 }{ -4-1 }=\frac{ -15 }{ -5 }=3\] Sorry this is what I meant. You still get the same answer.